Find the range of:


f(x)=x2-3x+2,0≤x≤4

Grade:12th Pass

2 Answers

Chetan Mandayam Nayakar
12 years ago

x^3 -3x+2=(x-2)(x-1)=(x-(3/2))2-(1/4), minimum=-1/4 when x=3/2, maximun=6 when x=4

range=[-1/4,6]

Ashwin Muralidharan IIT Madras
12 years ago

Hi Meanka,

 

The solution is in the scanned copy:

Hope that helps.

 

All the best.

Regards,

Ashwin (IIT Madras).